How many atoms do mercury vapour molecules consist of if if the density of mercury vapoue relative to air is 6.92?the average mass of air ia 29 g/ mole.

Grade:

1 Answers

Sunil Kumar FP
10 years ago
relative density=density of mercury/density of air=mass of 1 mole of hg/mass of 1 mole of air
=mass of 1 mole of mercury=6.92*28=200.68 g/mole=molecular mass of Hg
no of atom=avagadros no=6.023*10^23

thanks and regards
sunil kr
askIItian faculty